Thedreaded pathogenic bacterium J. grocious makes a toxin called BalD. There are two allelic forms of balD, with the balD-1 allele encoding the more potent form of the toxin, and balD-2 allele encoding the less potent form of the toxin. Every person in Champaign-Urbana is infected with one of two possible strains of J. grocious, which differ only in whether the strain carries balD-1 allele or balD-2 allele. If the allelic frequency of J. grocious strains carrying the balD-1 allele is 0.8 in Champaign (population = 80,000), and 0.6 in (Urbana = 50,000), then what is the allelic frequency of J. grocious strains carrying the less toxic balD allele within the population of the combined cities of Champaign-Urbana?

Options:

a. 0.72
b. 0.14
c. 0.28
d. 0.86
e. 0.42
Answer #1

Ans: As described in the question the toxin BalD has two allelic forms balD-1 and bal-D2. The allelic frequency of balD-1 in Champaign is 0.8 and 0.6 in Urbana and we have to find the allelic frequency of balD-2 within the population of the combined cities. According to Hardy Weinberg equation p+q=1 where p is the frequency of allele A and q is the frequency of allele B, so for Champaign the allelic frequency for less toxic strain is 0.2. Now allelic frequency is basically no. of organisms carrying that allele divided by total number of organisms so putting the values of each components we get the number of organisms carrying less toxic strain in Champaign= 0.2*80,000= 16000. Similarly for Urbana allelic frequency for less toxic allele is 0.4 and number of organisms carrying that allele is 0.4*50,000= 20,000.

Now when we combine the two population we get the total population 130,000 out of which 36,000 contains less toxic allele so its allelic frequency is equal to 36,000/130,000= 0.27692 which is approximately equal to 0.28 so the correct answer is option c.